the ratio of two number is 5:9 if their difference is 64. Find the bigger number.
Answer by macston(5194)

Let L=larger number and S=smaller number
5L=9S
L-S=64: L=S+64 Solve for L and substitute
5(S+64)=9S
5S+320=9S Subtract 5S from each side
320=4S Divide each side by 4
80=S The smaller number is 80
L-S=64
L-80=64 Add 80 to each side
L=144 ANSWER The larger number is 144
CHECK:
5L=9S
5(144)=9(80)
720=720